OpenStudy (anonymous):

Solve the following system of linear equations by addition: 8x-9y=1 1.6x-1.6y=16

OpenStudy (anonymous):

We have the equations- 8x-9y=1 -1st eq 1.6x-1.6y=16 -2nd eq Dividing the second equation throughout by -1.6 we have - -x+y=-10 -3rd eq Now multiply the 3rd eq by 8, we have -8x+8y=-80 -4th eq Now add the 4th and 1st equations, 8x -9y=1 -8x+8y=-80 Therefore -8x and +8x get cancelled and we have, -y=-79 Therefore y=79 and since -x+y=-10 substituting the value we have- -x +79=-10 Therefore x=-89 Answer, x=-89, y=79 !!!!